i need help with this problem:
n-3 n-4
---=---
n+5 n+3

Answer by checkley77(12569)

n-3 n-4
---=---
n+5 n+3 cross multiply:
(n+5)(n-4)=(n-3)(n+3)
n^2+n-20=n^2-9
n^2-n^2+n=20-9
n=11 answer.